OpenStudy (anonymous):

Eliminate the parameter. x = 3 cos t, y = 3 sin t

OpenStudy (anonymous):

are u miley cyrus..... :O

OpenStudy (anonymous):

ahhhhhh

OpenStudy (campbell_st):

take the 1st equation and make t the subject \[\frac{x}{3} = \cos(t).... then.... t = \cos^{-1}(\frac{x}{3})\] now just substitute this into the 2nd equation for t.

OpenStudy (anonymous):

how do i do that

OpenStudy (campbell_st):

like this \[y = 3\sin(\cos^{-1} (\frac{x}{3}))\]

OpenStudy (anonymous):

2.828

OpenStudy (campbell_st):

there is no numeric answer that I am aware or as you don't have a value for x.

OpenStudy (anonymous):

are u miley cyrus!!! :)

OpenStudy (anonymous):

oh...do i need to find the value of x